4 gm of MgCO3 on strong heating gives 896 ml of co2 at stop .what is the percentage purity of the sample?

Grade:12th pass

1 Answers

Himanshu Rathour
5 years ago
Reaction involved 
MgCO3 heat> MgO+CO2
1mole                  1mole  1mole
Mole of MgCO3 taken=4/85=0.047
1 mole produces 1 moles CO2
0.047 mole produces 0.047 mole CO2
But it is given CO2 produced=896/22400=0.04moles
% purity=    (0.047-0.04)x100
                        0.047
=14.89%